# Beam Splitter Physics and Testing Bell Inequalities with Phase Retarders

### v2

@@ Abstract
The role of Beam Splitters (BS) is crucial for quantum physics as it reveals the statistical behavior of quantum mechanical particles (quantons) and some of the fundamental quantum phenomena such as quantum superposition and randomness. Here, we investigate the use of BS and Phase Retarders (P) in combination and emphasize the importance of BS-P-BS systems for emerging quantum technologies. We demonstrate that the detection probabilities are equivalent to the spin measurement probabilities in analog setups. Then, by extending the discussion to spatially correlated two-quanton systems, we construct a suitable basis for analysis of other quantum mechanical processes that seem to have common origins. Last, we propose a new experimental setup for testing Bell-CHSH inequalities with spatially correlated two-quanton BS-P-BS systems by introducing topological phases.
preprint: APS/123-QED
Beam splitters (BS) are the most crucial components of theoretical or experimental studies that demonstrate the behaviors of photons as particles and the existence of physical properties like non-locality or indistinguishability. The main purpose of using BS appears to be quite straightforward from the classical perspective; however, its significance in quantum physics is far more profound [1]. Considering the input and output ports of the BS as quantum states, one can simulate quantum transport [2], create entangled quantum states [3], or build a universal quantum computer [4].
On the other hand, special properties emerge when BSs are utilized in combination with phase retarders (P). BS with P systems (BS-P-BS) reveal strong connections between spin measurement and detection of the quantum mechanical particles (quantons). Accordingly, the probabilities of being spin up or down in an arbitrary direction are equivalent to the probabilities of quantons being detected by detectors. Also, the phase of the retarder corresponds to the angle of the direction in which the spin of quantons is measured.
The importance of using BS-P-BS is far more evident in systems consisting of two spatially correlated (or analogously spin entangled) quantons. Such a two-quanton setup can be designed in two different configurations as either quantons produced from the same source reach the BSs after Ps or into the BS-P-BS system in an EPR-Bohm-type setup [5-6]. Since the BS-P-BS structure contains closed trajectories for each quanton, unlike the other configuration, it allows the study of other quantum mechanical/physical effects that seem to have common origins. Hence, the effects of topological phases on spatially correlated systems can be investigated by introducing electromagnetic field sources at the singular region where the quantons are not allowed to move. In this way, the detection probabilities can be compared with the results in a recent study for spin-entangled quantons in the framework of entirely relativistic quantum mechanics [7]. Moreover, at this stage, it is possible to define an observable in terms of the detection probabilities and to find the expectation values and correlation functions. Thus, the EPR-Bohm-type BS-P-BS setup can be used as a new experimental test method for Bell-CHSH inequalities (BI) ensuring a limit expected to be satisfied by deterministic local-hidden-variable-models (LHMV) [8-9].
@@
The first indication of this was provided in a different sense by Silverman [10]. In this work, Silverman aimed to reveal the amplitudes for joint-detection probabilities of quantons arrival at two detectors without P, and the effects of spin and Aharonov-Bohm (AB) phase on these probabilities by taking into account reflection (r) and transmission (t) coefficients of BS [11]. He proposed various EPR-Bohm-type hybrid setups consisting of a pair of spatially correlated quantons generated from the same source to obtain the correlation functions that can be controlled by the magnetic flux sourced by solenoids placed at singular points in closed trajectories formed by mirrors and spin-selective BS (SBS). The outcomes of Silverman’s work are quite inspiring in many respects. However, the AB phase is a spin-independent effect, and the proposed setup only allows us to find the detection probabilities without any spin dependency [7]. Hence, the relative phase information does not appear in the expectation value expressions.
In this letter, we demonstrate that the joint-detection probabilities are equivalent to the joint-spin measurement probabilities for all kinds of quantons (bosons, fermions) by introducing BS-P-BS systems. Accordingly, we reveal that the angles of the arbitrary directions of the spin measurements can be arranged by the phases of the retarders. Furthermore, we modify [12] with the Aharonov-Casher phase that arises as a result of the motion of the magnetic dipole moment carrier neutrons around the linear charge distributions by illustrating the effects of topological phases on two-quanton correlated systems. Last, we propose a new gedanken experimental setup for testing BI in BS-P-BS systems.
@@
Beam Splitters.—Consider a symmetric and lossless beam splitter (BS) with two ports 00 and $1$ for incoming, and two ports $D_{0}$ and $D_{1}$ for outgoing quantons (See Fig. 1.a). The basis $(1\thinspace,0)^{T}\equiv\left|0\right\rangle$ and $(0\thinspace,1)^{T}\equiv\left|1\right\rangle$ for ports 0 and 1 can be used to describe the state $\left|\psi^{i}_{in}\right\rangle$ of incoming quantons entering such a BS through the port $i$ $(i=0,1)$. Similarly, to represent the output states one can choose the basis $(1\thinspace,0)^{T}$ and $(0\thinspace,1)^{T}$ for $\left|D_{0}\right\rangle$ and $\left|D_{1}\right\rangle$. Hence, the BS performs a linear transformation that converts $\left|\psi^{i}_{in}\right\rangle$ to $\left|\psi_{out}\right\rangle$. By denoting purely real $(r_{i}=R_{i})$ and purely imaginary $(t_{i}=iT_{i})$ as reflection and transmission amplitudes of the BS respectively (Namely, for incoming quantons through port 0 as $r_{0}$ and $t_{0}$, and through port 1 as $r_{1}$ and $t_{1}$), $\left|r_{0}\right|=\left|r_{1}\right|=\left|t_{0}\right|=\left|t_{1}\right|$; $\left|r_{i}\right|^{2}+\left|t_{i}\right|^{2}=1$; $r_{0}^{*}t_{1}+t_{0}^{*}r_{1}=0$ are provided. Therefore the operation of the BS can be given by a unitary matrix [1],
@@
Figure 2: A proposal for the use of BS-P for spatially correlated two-quanton systems generated from the same source
On the other hand, some significant physical characteristics appear when a second BS is added with a phase retarder, and the detection probabilities of quantons become dependent on the phase of the retarder (see Fig. 1.b). Furthermore, this phase corresponds to the angle of the arbitrary direction on which spin measurements are performed to find the probabilities (up or down). According to the existing configuration, wlog retarder operator ${P}$ can be expressed as,
@@
In the case of single incoming quanton through port $i$, system evolves into the state $\left|\psi^{i}_{in}\right\rangle\rightarrow\left|\psi_{out}\right\rangle={[BS_{2}][P][BS_{1}]}\left|\psi^{i}_{in}\right\rangle$. After $\mathrm{BS_{1}}$ the initial state of the system transforms into $\left|\psi^{i}_{in}\right\rangle\rightarrow\frac{1}{\sqrt{2}}\left[\left|\psi_{0}^{{}^{\prime}}\right\rangle+i\left|\psi_{1}^{{}^{\prime}}\right\rangle\right]$, and gains an additional phase factor via the retarder $\left|\psi_{0}^{{}^{\prime}}\right\rangle\rightarrow e^{i\vartheta}\left|\psi_{0}^{{}^{\prime}}\right\rangle=\frac{e^{i\vartheta}}{\sqrt{2}}\left[i\left|{D_{0}}\right\rangle+\left|{D_{1}}\right\rangle\right]$. Consequently, the final state of the system can be given in a compact representation,
@@
The term $(ie^{i\frac{\vartheta}{2}})$ in (3) is the common phase factor and can be neglected for both two equations. These results are valid for all kinds of quantons (bosons, fermions) and just related with the spatial part of the particle’s wavefunctions and one can easily calculate the probabilities using (3).
Now, we extend the problem to investigate possible uses of BS in correlated systems in two scenarios. First, let’s assume that two spatially-correlated (or simply momentum-correlated) quantons produced from the same source, and follow either orange or blue lines as in the Fig. 2. In this scenario, the initial state of the system can be considered as,
@@
and one can use (2) as a retarder operator with the selection of the phase $\vartheta_{L(R)}$ consistent with the path. The correlated quantons reach BSs after passing through reflecting mirrors and retarders on both sides. Thus, the state of the system (4) transformed into the state $\left|\psi_{out}\right\rangle=(BS_{L}\otimes BS_{R})(P_{L}\otimes P_{R})\left|\psi_{in}\right\rangle$. It can be written in an explicit form by neglecting the overall phase factor,
@@
$-e^{i\vartheta}[\thinspace i|{D_{0}}^{\prime}\rangle+|{D_{1}}^{\prime}\rangle]\otimes[\thinspace|{D_{0}}\rangle+i|{D_{1}}\rangle])$
@@
$A(D^{\prime}_{1},D_{0})\equiv\left\langle{D^{\prime}_{1}}{D_{0}}\right|\left.\psi_{out}\right\rangle=-\frac{1}{2\sqrt{2}}(1+e^{i\vartheta})$ (7)
@@
$P(D_{1}^{{}^{\prime}},D_{0})\equiv\left\|A(D_{1}^{{}^{\prime}},D_{0})\right\|^{2}=\frac{1}{2}\sin^{2}\left(\frac{\vartheta_{L}-\vartheta_{R}}{2}\right).$ (9)
These results are related with the joint-spin measurement probabilities obtained in EPR-Bohm type experimental setups using spin-correlated particles [5-6], with a phase shift $\pi/2$ as Degiorgio’s relation due to BS [13]. It is also clear that the probabilities depend on the phase of the retarder.
Second, a new BS-P-BS system for a correlated quanton pair can be proposed to establish a comprehensive analogy between spin measurements and detection processes. Accordingly, spatially-correlated quantons produced in the state (4) are subjected to retarders placed on one of the arms after passing the first BSs. Then, they move through the second BSs and reach the detectors (See Fig 3). In this new configuration, the final state of the system is obtained as $\left|\psi_{out}\right\rangle=(BS_{L_{1}}\otimes BS_{R_{1}})(P_{L}\otimes P_{R})(BS_{L_{2}}\otimes BS_{R_{2}})\left|\psi_{in}\right\rangle$, or clearly,
@@
Thus, one can easily find the joint-detection probabilities for quantons at $D_{0}$ and $D_{1}$ (or $D_{1}^{\prime}$):
$P(D_{0}^{{}^{\prime}},D_{0})=\frac{1}{2}\sin^{2}\left(\frac{\vartheta_{L}-\vartheta_{R}}{2}\right)\equiv P(\alpha\Uparrow_{\tilde{n}},\beta\Uparrow_{\tilde{n}})$ (11)
$P(D_{1}^{{}^{\prime}},D_{0})=\frac{1}{2}\cos^{2}{\left(\frac{\vartheta_{L}-\vartheta_{R}}{2}\right)}\equiv P(\alpha\Downarrow_{\tilde{n}},\beta\Uparrow_{\tilde{n}})$ (12)
The above results are well known expressions for probabilities of joint-spin measurements in entangled quantum systems [7]. The phase of the retarder can be thought of as the angle of the arbitrarily chosen direction $\hat{\bm{n}}$ in which the spin is measured in an analog experimental setup. In this regard, (9) clearly indicates the connections between these processes without any Degiorgio phase shift. Accordingly, the probabilities of spin up (or down) in $\hat{\bm{n}}$ direction are equivalent to the probabilities of detected by a detector placed up (or down) for the corresponding phase of the retarder to the angle of n with horizontal.
Figure 3: The use of the BS-P-BS system based on detection probabilities as an analog to spin measurement probabilities for spatially-correlated two-quanton systems and a proposal for a gedanken experimental setup for testing BI.
Moreover, the proposed setup can be used to test BI which provides a limit expected to be satisfied by the deterministic LHVM [5,9]. The violation of BI is meaningful for demonstrating the existence of non-local properties of quantum mechanics. Therefore, in EPR-Bohm-type experimental setups, one tries to find the relevant correlation function S as a combination of expectation value expressions based on different spin-measurement probabilities. In general, to obtain the correlation function S in an entangled two spin-1/2 particles in a singlet state
@@
one usually starts with decomposing it into the instantaneous eigenstates of the system in any arbitrarily chosen direction $\bm{\hat{n}}$, then observes the joint-spin measurement probabilities of particles in the directions that make different $\alpha$ and $\beta$ angles with $\bm{\hat{n}}$ on both sides. In our configuration, the spin measurement probabilities is replaced by the detection probabilities.
Figure 4: A proposal for testing the possible effects of the AB phase on entangled quantum states.
As an illustration (See Fig 4), joint-probabilities for the particle on the left finding spin up on the direction making an angle $(\vartheta_{L}\equiv\alpha)$ and for the particle on the right finding on the direction making an angle $(\vartheta_{R}\equiv\beta)$ spin up $P\left(\bm{\alpha}\Uparrow_{\bm{\hat{n}}},\bm{\beta}\Uparrow_{\bm{\hat{n}}}\right)$ or down $P\left(\bm{\alpha}\Uparrow_{\bm{\hat{n}}},\bm{\beta}\Downarrow_{\bm{\hat{n}}}\right)$ with the direction $\bm{\hat{n}}$ are exactly same with the joint-spin probabilities $P(D_{0}^{{}^{\prime}},D_{0})$ and $P(D^{\prime}_{1},D_{0})$ found in (9). By introducing the observable $\mathcal{O}$,
$\mathcal{O}^{l}(\vartheta_{L})=P_{D_{0}^{{}^{\prime}}}^{l}(\vartheta_{L})-P_{D_{1}^{{}^{\prime}}}^{l}(\vartheta_{L})$ (14)
and similarly $\mathcal{O}^{r}(\vartheta_{R})$ for the angle $\vartheta_{R}$, one calculate the expectation value of joint measurements,
$E(\vartheta_{L},\vartheta_{R})$ $=$ $\left\langle\psi\left(\tau\right)\right|\mathcal{O}^{l}(\vartheta_{L})\otimes\mathcal{O}^{r}(\vartheta_{R})\left|\psi\left(\tau\right)\right\rangle$ (15)
$=$ $-\cos{(\vartheta_{L}-\vartheta_{R})}$
@@
With the certain chosen angles for the correlation function $S(0,\frac{\pi}{4},\frac{3\pi}{4},\frac{\pi}{2})=2\sqrt{2}>2$, BI is maximally violated by the quantum mechanical expectation values. Historically, violations of BI have been demonstrated in numerous experimental studies using neutral kaons or similar large systems [14-20]. However, the proposed setup is quite simple compared to its counterparts. It is easy to implement in a laboratory environment with all kinds of quantons (bosons or fermions). It has fully closed trajectories on both sides (in contrast to the preceding one), and it enables the investigation of other physical/quantum mechanical systems as topological phases, which is particularly important for experimental studies and emerging quantum technologies. Besides, it is also very noticeable to reach that conclusion based on the probabilities of being detected, without any spin information. In order to use a spin-entangled particle pair instead, it would be sufficient to utilize spin-selective BSs (SBSs) (which reflect all spin down and pass all spin up particles).
Figure 5: (a) Schematic representation of the AC phase in 2+1 dimensions resulting from the motion of magnetic dipole moment carrier chargeless particles in a closed trajectory around a linear charge distribution. (b) A proposal for a gedanken experimental setup to investigate the effects of topological phases on entangled quantum states and to observe the violations of BI using physical quantities in the singular region to test the non-local properties of quantum mechanics.
Testing Bell inequalities with phase retarders.—To provide an illustration of the use of BS detailed in the previous section, we reconsider proposed setups by introducing a topological phase. Hence, we observe the contribution of particle spins to the process and make the function S given by (13) for BI dependent on the topological phase. Topological phases are quantum mechanical phenomena that occur in multiply-connected region under the influence of vector potential or vector potential-like physical quantities, without classical Lorentz forces, independent of the geometry of the space and the speed of the particles [11-12, 21-25]. It is also invariant under spatial translations (traditionally in the z-direction), and essentially arises in two-dimensional space. The first and most well-known example of topological phases is the AB phase, which is based on the control of the interference pattern via an infinitely long and thin solenoid placed beyond the reach of moving electrons in a double-slit experiment. However, the interaction Hamiltonian describing the dynamics of the AB system and the transformed wavefunction are given by,
where e is the electron charge, $\bm{\alpha}$ is the Pauli matrices, and A is the vector potential in 2+1 dimensions (for detailed information see ref: [7]). The interaction (between the particles and vector potential sourced by the solenoid), and the phase are independent of the spin orientation of the particles as is seen from (14). Thus, the relative phase in spin-entangled systems disappears in the expectation value expressions for the AB case contrarily done in Silverman’s study [10]. The results are reduced to the equations (8) and (9) even in the presence of the solenoid in proposed setups.
Instead, it is appropriate to use the AC phase arising in a stationary reference frame of AB electrons in the presence of unpolarized magnetic dipoles [12]. The AC phase is based on controlling the interference pattern with the linear charge distribution in the singular region in a double-slit experiment with neutrons carrying magnetic dipole moments without the effects of the classical forces. The AC Hamiltonian describing the interaction, and the transformed wave function are as follows,
$\begin{matrix}\Delta H_{AC}=-s\bm{\mu}\cdot\bm{\tilde{E}}\\ \psi^{\prime}=e^{-is\mu\lambda}\psi_{0}\end{matrix}$ (18)

### v3

# Testing Bell-CHSH Inequalities Using topological Aharonov-Casher and He-McKellar-Wilkens Phases

@@ Abstract
The effects of Aharonov-Casher (AC) and He-McKellar-Wilkens (HMW) phases on entangled spin-1/2 quantum systems are investigated. We use linear charge distributions positioned at the center of resulting closed orbits, capitalizing on Mach Zender-type interferometers modified with phase retarders to unveil the topological effects. We analyze how AC-HMW phases influence the Bell angles and maximal violation of Bell-CHSH inequalities (BI) without any classical interaction. We incorporate the spin and path of particles in the interferometers as observables to test noncontextual hidden variable theories against quantum mechanics, leveraging the non-local features of AC-HMW effects. Furthermore, we discuss potential implementations of our scheme in physical systems.
keywords:
Bell-CHSH Inequalities , Aharonov-Casher Phase , He-McKellar-Wilkens Phase , Noncontextuality
[…]
1 Introduction
Testing Bell-CHSH inequalities (BI) which impose constraints on the correlations observed among distinct components within multipartite systems within the framework of local realism is crucial for serving as experimental validations of the foundational principles of quantum mechanics, particularly the concept of quantum entanglement  [1; 2; 3]. BI has been tested in diverse quantum systems such as photons and atoms in quantum optics, high-energy physics, neutrals, charmonium decays, neutrino oscillations  [4; 5; 6; 7; 8; 9; 10; 11; 12; 13; 14; 15; 16], and the experiments indicate the existence of non-local features of quantum mechanics.
On the other hand, the topological properties of physical systems are crucial for quantum communication and quantum computation applications with their non-local features. One of the best-known and earliest examples of this is the Aharonov-Bohm (AB) physical/quantum mechanical process  [17; 18; 19]. AB-type effects are explained by the introduction of the vector potential or vector potential-like physical quantities as complex phase factors into the wave functions of particles moving along the closed trajectories around singular regions created by electromagnetic field sources without any classical interactions. Also, AB-type effects are invariant under spatial translations through the z direction (traditionally) due to the symmetries of the problems, and there are duality and identity relationships between phases particularly appearing in two dimensions  [20; 21]. Accordingly, while AB and its fully dual-AB (DAB) effects are independent of the spin orientations of moving particles, Aharonov-Casher (AC) and its fully dual He-McKellar-Wilkens (HMW) effects demonstrate dependence on spin polarizations  [22; 23; 24; 25; 26; 27]. Between AB and AC (similarly between HMW and DAB) one can speak of an identity rather than a duality that arises in the static reference frames of charges, especially if the dipoles are polarized. In this regard, AC-HMW effects serve as practical tools for testing nonlocality and noncontextuality in comparison to quantum mechanics within EPR-Bohm-type experiments  [28; 29; 30; 31].
@@ 1 Introduction
In this work, we propose a scheme to observe the effects of topological phases on entangled quantum states by exploiting the non-local structure of the AC phase and to test of incompatibility of noncontextuality with quantum mechanics, which is a restrictive demand for a theory, via BI without any classical interaction or adiabaticity condition. For this purpose, we use a pair of Mach Zender-type interferometers modified with phase retarders to be analogous to spin measurement angles in arbitrary directions (BS-P-BS). We place the electromagnetic field source (linear charge distributions) in the closed regions between the legs of the interferometers, which creates the topological phase contributions to the wave functions of the moving particles (dipoles). We demonstrate that the correlation function becomes dependent on the topological phase by obtaining the detection probabilities (or spin measurement probabilities) for fixed phases of retarders (or Bell angles) maximally violated by quantum mechanical expectation values after passing through mutual BS-P-BS systems.
This letter is organized as follows. In the second section, we provide the background of the study as a preliminary preparation. We review the quantum mechanical use of Beam-Splitters (BS) with phase retarders (P) and consider the possible scenarios for BI testing for all kinds of quantum mechanical particles (quantons), not only photons, by the existing notation. In the third section, we first modify the proposed scenarios with the contribution of the AC phase. Emphasizing the importance of spin dependences of moving particles, we obtain the correlation function for testing BI in the BS-P-BS system that becomes AC-HMW phase dependent. Lastly, we discuss the possible implementations of our scheme in physical systems and a generalization that can be used for other studies in the literature.
2 Notation and Background
Beam splitters (BS) are the most crucial components of theoretical or experimental studies that demonstrate the behaviors of photons as particles and the existence of physical properties like non-locality or indistinguishability. The main purpose of using BS appears to be straightforward from the classical perspective; however, its significance in quantum physics is far more profound  [32]. Moreover, special properties emerge when BSs are utilized in combination with phase retarders (P) in Mach-Zender interferometers (BS-P-BS) revealing strong connections between spin measurement and detection of the quantum mechanical particles (quantons). Accordingly, the probabilities of being spin up or down in an arbitrary direction are equivalent to the probabilities of quantons being detected by detectors. The phase of the retarder corresponds to the angle of the direction in which the spin of quantons is measured. In this way, they can be easily used in systems consisting of two spatially correlated (or similarly spin-entangled) quantons in EPR-Bohm-type experiments  [33]. Besides, the BS-P-BS structure that contains closed trajectories for each quanton allows us to study other quantum mechanical/physical effects that seem to have common origins. Hence, the effects of topological phases on spatially correlated systems can be investigated by introducing electromagnetic field sources at the singular region where the quantons are not allowed to move. In this way, the detection probabilities can be achieved for spin-entangled quantons to find the expectation values and correlation functions. Thus, it can be used for testing BI without photons  [2; 3].
@@ 2 Notation and Background
Lets consider a symmetric and lossless BS with two ports 00 and $1$ for incoming, and two ports $D_{0}$ and $D_{1}$ for outgoing quantons (See Fig. 1.a). The basis $(1\thinspace,0)^{T}\equiv\left|0\right\rangle$ and $(0\thinspace,1)^{T}\equiv\left|1\right\rangle$ for ports 0 and 1 can be used to describe the state $\left|\psi^{i}_{in}\right\rangle$ of incoming quantons entering such a BS through the port $i$ $(i=0,1)$. Similarly, to represent the output states one can choose the basis $(1\thinspace,0)^{T}$ and $(0\thinspace,1)^{T}$ for $\left|D_{0}\right\rangle$ and $\left|D_{1}\right\rangle$. Hence, the BS performs a linear transformation that converts $\left|\psi^{i}_{in}\right\rangle$ to $\left|\psi_{out}\right\rangle$. By denoting purely real $(r_{i}=R_{i})$ and purely imaginary $(t_{i}=iT_{i})$ as reflection and transmission amplitudes of the BS respectively (Namely, for incoming quantons through port 0 as $r_{0}$ and $t_{0}$, and through port 1 as $r_{1}$ and $t_{1}$), $\left|r_{0}\right|=\left|r_{1}\right|=\left|t_{0}\right|=\left|t_{1}\right|$; $\left|r_{i}\right|^{2}+\left|t_{i}\right|^{2}=1$; $r_{0}^{*}t_{1}+t_{0}^{*}r_{1}=0$ are provided. Therefore the operation of the BS can be given by a unitary matrix,
@@
Figure 2: A scheme for the use of BS-P for spatially correlated two-quanton systems generated from the same source
@@
On the other hand, some significant characteristics appear when a second BS is added with a phase retarder. The detection probabilities of quantons become dependent on the phase of the retarder (See Fig. 1.b). Furthermore, this phase corresponds to the angle of the arbitrary direction on which spin measurements are performed to find the probabilities (up or down). According to the existing configuration, wlog, the retarder operator ${P}$ can be expressed as,
@@
In the case of single incoming quanton through port $i$, system evolves into the state $\left|\psi^{i}_{in}\right\rangle\rightarrow\left|\psi_{out}\right\rangle={[BS_{2}][P][BS_{1}]}\left|\psi^{i}_{in}\right\rangle=\frac{e^{i\vartheta}}{\sqrt{2}}\left[i\left|{D_{0}}\right\rangle+\left|{D_{1}}\right\rangle\right]$. Consequently, the final state of the system can be given in a compact representation,
@@
The term $(ie^{i\frac{\vartheta}{2}})$ is the common phase factor and can be neglected for both two equations. These results are valid for all kinds of quantons (bosons, fermions) and just related to the spatial part of the particle’s wavefunctions.
Now, we extend the problem in correlated systems in two scenarios. First, let’s assume that two spatially correlated quantons are produced from the same source, and follow either orange or blue lines as in Fig. 2. In this scenario, the initial state of the system can be considered as,
@@
and one can use  (2) as retarder operator with the selection of the phase $\vartheta_{L(R)}$ consistent with the path. The correlated quantons reach BSs after passing through reflecting mirrors and retarders on both sides. Thus, the state of the system  (4) transformed into the state $\left|\psi_{out}\right\rangle=(BS_{L}\otimes BS_{R})(P_{L}\otimes P_{R})\left|\psi_{in}\right\rangle$. It can be written in an explicit form by neglecting the overall phase factor,
@@
$\quad-e^{i\vartheta}[\thinspace i|{D_{0}}^{\prime}\rangle+|{D_{1}}^{\prime}\rangle]\otimes[\thinspace|{D_{0}}\rangle+i|{D_{1}}\rangle])$
@@
$A(D^{\prime}_{1},D_{0})$ $\equiv\left\langle{D^{\prime}_{1}}{D_{0}}\right|\left.\psi_{out}\right\rangle=-\frac{1}{2\sqrt{2}}(1+e^{i\vartheta})$
@@
$P(D_{1}^{{}^{\prime}},D_{0})$ $\equiv\left\|A(D_{1}^{{}^{\prime}},D_{0})\right\|^{2}=\frac{1}{2}\sin^{2}\left(\frac{\vartheta_{L}-\vartheta_{R}}{2}\right).$
These results are related to the joint-spin measurement probabilities obtained in EPR-Bohm type experimental setups using spin-correlated particles, with a phase shift $\pi/2$ as Degiorgio’s relation due to BS  [34]. It is also clear that the probabilities depend on the phase of the retarder.
Second, BS-P-BS system can be proposed to establish an analogy between spin measurements and detection processes. Accordingly, spatially correlated quantons produced in the state  (4) are subjected to retarders placed on one of the arms after passing the first BSs. Then, they move through the second BSs and reach the detectors (See Fig. 3). In this new configuration, the final state of the system is obtained as $\left|\psi_{out}\right\rangle=(BS_{L_{1}}\otimes BS_{R_{1}})(P_{L}\otimes P_{R})(BS_{L_{2}}\otimes BS_{R_{2}})\left|\psi_{in}\right\rangle$, or clearly,
@@
Thus, one can find the joint-detection probabilities for quantons at $D_{0}$ and $D_{0}^{\prime}$ (or $D_{1}^{\prime}$):
$P(D_{0}^{{}^{\prime}},D_{0})=\frac{1}{2}\sin^{2}\left(\frac{\vartheta_{L}-\vartheta_{R}}{2}\right)\equiv P(\vartheta_{L}\Uparrow_{\tilde{n}},\vartheta_{R}\Uparrow_{\tilde{n}})$
$P(D_{1}^{{}^{\prime}},D_{0})=\frac{1}{2}\cos^{2}{\left(\frac{\vartheta_{L}-\vartheta_{R}}{2}\right)}\equiv P(\vartheta_{L}\Downarrow_{\tilde{n}},\vartheta_{R}\Uparrow_{\tilde{n}})$ (9)
The above results are well-known expressions in entangled quantum systems  [21]. The retarder phase corresponds to the spin measurement angle $\hat{\bm{n}}$. Equation  (2) clearly connects these processes without any Degiorgio phase shift. Accordingly, the probabilities of spin up (or down) in $\hat{\bm{n}}$ direction are equivalent to the probabilities detected by a detector placed up (or down) for the corresponding phase of the retarder to the angle of n with horizontal. Moreover, this setup can be used for testing BI, crucial for demonstrating non-local properties of quantum mechanics. Therefore, in EPR-Bohm-type experimental setups, one tries to find the relevant correlation function S as a combination of expectation value expressions based on different spin-measurement probabilities. In general, to obtain the correlation function S in entangled two spin-1/2 particles in a singlet state,
Figure 3: The use of the BS-P-BS system based on detection probabilities as an analog to spin measurement probabilities for spatially correlated two-quanton systems and a scheme for a gedanken experimental setup for testing BI.
@@
One can start with decomposing it into the instantaneous eigenstates of the system in any arbitrarily chosen direction $\bm{\hat{n}}$, then observe the joint-spin measurement probabilities of particles in the directions that make different $\vartheta_{L}$ and $\vartheta_{R}$ angles with $\bm{\hat{n}}$ on both sides. Here, the spin measurement probabilities are replaced by the detection probabilities. In other words, $P\left(\bm{\vartheta_{L}}\Uparrow_{\bm{\hat{n}}},\bm{\vartheta_{R}}\Uparrow_{\bm{\hat{n}}}\right)$ and $P\left(\bm{\vartheta_{L}}\Uparrow_{\bm{\hat{n}}},\bm{\vartheta_{R}}\Downarrow_{\bm{\hat{n}}}\right)$ are exactly same with $P(D_{0}^{{}^{\prime}},D_{0})$ and $P(D^{\prime}_{1},D_{0})$. By introducing the observable $\mathcal{O}$,
$\mathcal{O}^{l(r)}(\vartheta_{L}(R))=P_{D_{0}^{{}^{\prime}}}^{l(r)}(\vartheta_{L}(R))-P_{D_{1}^{{}^{\prime}}}^{l(r)}(\vartheta_{L}(R))$ (11)
one can calculate the expectation values,
$E(\vartheta_{L},\vartheta_{R})=\left\langle\psi\left(\tau\right)\right|\mathcal{O}^{l}(\vartheta_{L})\otimes\mathcal{O}^{r}(\vartheta_{R})\left|\psi\left(\tau\right)\right\rangle=-\cos{(\vartheta_{L}-\vartheta_{R})}$ (12)
@@
With the certain chosen angles for the correlation function $S(0,\frac{\pi}{4},\frac{3\pi}{4},\frac{\pi}{2})=2\sqrt{2}>2$, BI is maximally violated by the quantum mechanical expectation values.
3 Testing Bell-CHSH inequalities using topological phases
Topological phases are quantum mechanical phenomena that occur in multiply-connected regions under the influence of vector potential or vector potential-like physical quantities, without classical Lorentz forces, independent of the geometry of the space and the speed of the particles  [17; 18; 19]. It is also invariant under spatial translations (traditionally in the z-direction) and essentially arises in two-dimensional space. The first and most well-known example of topological phases is the AB phase, which is based on the control of the interference pattern via an infinitely long and thin solenoid placed beyond the reach of moving electrons in a double-slit experiment. Subsequently, consequential studies have been performed and experimentally observed, including duality and identity relations with the AB system  [20; 21; 22; 23; 24; 25; 26; 27; 35; 36; 37; 38; 39; 40; 41; 42; 43; 44; 45; 46; 47; 48]. At the forefront of these is the AC effect arises from the motion of chargeless neutrons carrying a magnetic dipole in closed trajectories around an electrically linear charge distribution  [22; 23]. Similarly, the HMW effect occurs due to the motion of chargeless particles carrying an electric dipole around a linear charge distribution acting as a magnetic field source  [24; 25; 26]. The AC and HMW Hamiltonians describing the interaction and the transformed wave functions are,
$\begin{matrix}\Delta H_{AC}=-s\bm{\mu}\cdot\bm{\tilde{E}}&&&\Delta H_{HMW}=-s\bm{d}\cdot\bm{\tilde{B}}\\ \psi^{\prime}=e^{-is\mu\lambda_{E}}\psi_{0}&&&\psi^{\prime}=e^{-isd\lambda_{B}}\psi_{0}\end{matrix}$ (14)